OpenStudy (kmullis6):

Please help me.. I'm really confused. /:

OpenStudy (kmullis6):

I've already got part A... A = 9(x+23)/2 2A = 9(x+23) 2A/9 = (x+23) 2A/9 - 23 = x x = 2A/9 - 23

OpenStudy (kmullis6):

But I need help on the other 2.

OpenStudy (anonymous):

ok so take the equation that is had from part a, and instead of the variable, substitute it 108 then solve, u get it now?

OpenStudy (anonymous):

so plug 108 for a

OpenStudy (anonymous):

the solve for x

OpenStudy (kmullis6):

ok just a sec

OpenStudy (anonymous):

u need help?

OpenStudy (anonymous):

the equation is x=2a/9-23 so 2x108=216 216/9=24 24-23=1 x=1

OpenStudy (kmullis6):

So is it this? 108 = 9/2(x+23) 108 = 9(x+23)/2 216 = 9(x+23) 216/9 = x + 23 24 = x+23 24 - 23 = x 1 = x x = 1

OpenStudy (anonymous):

yes great!

OpenStudy (anonymous):

u need help on part c?

OpenStudy (kmullis6):

Yay! (: Yea, if you don't mind.

OpenStudy (anonymous):

ok for part c

OpenStudy (anonymous):

ok so we need to solve for n, this means we need to set an equation that solves for n so what we need to do is isolate the variable "n" so your steps are

OpenStudy (anonymous):

–np – 90 > 30 first add 90 to both sides -np>120 then you want to divide by p -n>120/p then finally u have to divide by -1 since there is an invisible 1 next to the p and we need to get rid of that negative n>-120/p is your final answer!

OpenStudy (kmullis6):

I was taught that if you divided by a negative, the sign was flipped. Am I confused about that?

OpenStudy (anonymous):

yes whenever u divide anything my a negative it becomes the oposite sign, youre correc about that!

OpenStudy (anonymous):

by*

OpenStudy (kmullis6):

So should the sign be flipped at the end in this?

OpenStudy (anonymous):

yes the answer is a negative fraction which is -120/p

OpenStudy (kmullis6):

Oh ok. Thank you so much for the help! (:

OpenStudy (anonymous):

np, wheneever u have a question, just tag me! id love to help

OpenStudy (kmullis6):

Thanks, I will if i need help.